OpenStudy (okdutchman7):

Please help. WILL MEDAL! Write the equation of the line equidistant from (8,4) and (-3,4) using the locus definition method. Write the equation in general form. @pooja195 @Kainui @sleepyjess @jim_thompson5910 @zepdrix @Photon336

OpenStudy (photon336):

@johnweldon1993

OpenStudy (okdutchman7):

I think I understand how to do most of it.

jimthompson5910 (jim_thompson5910):

I would use what you did earlier http://openstudy.com/users/okdutchman7#/updates/57313dfee4b071c0ef3cc3e8 use @mathmale 's method

OpenStudy (okdutchman7):

x = 5/2

OpenStudy (okdutchman7):

I just don't know how to convert it into general form.

OpenStudy (okdutchman7):

General form is: \[Ax + By + C = 0\]

OpenStudy (okdutchman7):

@Hero @robtobey

jimthompson5910 (jim_thompson5910):

well you can multiply both sides by 2, then subtract 5 from both sides x = 5/2 2x = 5 2x - 5 = 0 2x + 0y + (-5) = 0

jimthompson5910 (jim_thompson5910):

2x + 0y + (-5) = 0 is in Ax+By+C = 0 form A = 2 B = 0 C = -5

OpenStudy (okdutchman7):

So that would make it: 2x-5=0?

OpenStudy (okdutchman7):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes `2x-5 = 0`

OpenStudy (okdutchman7):

Thank You!

jimthompson5910 (jim_thompson5910):

you're welcome